OpenStudy (anonymous):

WILL GIVE A MEDAL. The slope of a line is 1, and the y-intercept is -1. What is the equation of the line written in slope-intercept form?

OpenStudy (texaschic101):

slope intercept form : y = mx + b The m is your slope and the b is your y intercept y = 1x + (-1) or it can be written like this : y = x - 1

OpenStudy (shadowlegendx):

@z053sme Slope-intercept form: y = mx + b y = any y point m = the slope x = any x point b = the y-intercept So lets input the given information The slope of a line is 1, and the y-intercept is -1. y = 1x - 1